1. Find the area under the standard normal curve that lies outside the interval between z= 0.57 and z= 1.82

2. Find the area under the standard normal curve to the left of z= 1.15

3. Find the area under the standard normal curve to the right of z= 2.04

4. Find the area under the standard normal curve that lies between z= =1.32 and z= 1.43
Answer #1

Solution,

Using standard normal table

1) P( Z < 0.57 OR Z > 1.82 )

= P(Z < 0.57 ) + P( Z > 1.82 )

= 0.7157 + 1 - P(Z < 1.82)

= 0.7157 + 1 - 0.9656

= 0.7157 + 0.0344

= 0.7501

2) P( Z < 1.15)

= 0.8749

3) P(Z > 2.04)

= 1 - P( Z < 2.04)

= 1 - 0.9793

= 0.0207

4) P( 1.32 < Z < 1.43)

= P( Z < 1.43) - P( Z < 1.32)

= 0.9236 - 0.9066

= 0.0170